The heights of adult men in America are normally distributed, with a mean of 69.6 inches and a standard deviation of 2.65 inches. The heights of adult women in America are also normally distributed, but with a mean of 64.3 inches and a standard deviation of 2.53 inches.
a) If a man is 6 feet 3 inches tall, what is his z-score (to two decimal places)?
z =

b) If a woman is 5 feet 11 inches tall, what is her z-score (to two decimal places)?
z =

c) Who is relatively taller?
-The 6 foot 3 inch American man
-The 5 foot 11 inch American woman

Answer by ewatrrr(24785)

Previously addressed:
Men Normal Distribution: µ = 69.1 and σ = 2.65
Women Normal Distribution: µ = 64.8 and σ = 2.59
z(man 75") = ( 75 - 69.1)/2.65 = 2.23
z(woman 71") = (71 - 64.8)/2.59 = 2.39*** The Woman.. relatively taller for a Woman
z-value larger (more unusual)